A manufacturing process is designed to produce bolts with a 0.5-inch diameter. Once each day, a random sample of 36 bolts istried 0.0025 too, did not work either
Answer #1

Solution :

Given that,

mean = \mu = 0.5

standard deviation = \sigma = 0.02

n = 36

\mu\bar x =  \mu = 0.5

\sigma\bar x = \sigma / \sqrt n = 0.02 / \sqrt 36 = 0.003

P(0.490 > \bar x OR \bar x > 0.510)  

P(\bar x < 0.490) = P((\bar x - \mu \bar x ) / \sigma \bar x < (0.490 - 0.5) / 0.003)

= P(z < -3.33)

Using z table

= 0.0004

P(\bar x > 0.510) = 1 - P(\bar x < 0.510)

= 1 - P[(\bar x - \mu \bar x ) / \sigma \bar x < (0.510 - 0.5) / 0.003]

= 1 - P(z < 3.33)

Using z table,    

= 1 - 0.9996

= 0.0004

= P(\bar x < 0.490 + P(\bar x > 0.510)

= 0.0004 + 0.0004

= 0.0008